# Making Offline RL Online: Collaborative World Models for Offline Visual Reinforcement Learning

### v3

@@ Abstract
Training offline reinforcement learning (RL) models using visual inputs poses two significant challenges, i.e., the overfitting problem in representation learning and the overestimation bias for expected future rewards. Recent work has attempted to alleviate the overestimation bias by encouraging conservative behaviors. This paper, in contrast, tries to build more flexible constraints for value estimation without impeding the exploration of potential advantages. The key idea is to leverage off-the-shelf RL simulators, which can be easily interacted with in an online manner, as the “test bed” for offline policies. To enable effective online-to-offline knowledge transfer, we introduce CoWorld, a model-based RL approach that mitigates cross-domain discrepancies in state and reward spaces. Experimental results demonstrate the effectiveness of CoWorld, outperforming existing RL approaches by large margins.
Machine Learning, ICML
@@ 1 Introduction
Learning control policies with visual observations can be challenging due to high interaction costs with the physical world. Offline reinforcement learning (RL) is a promising approach to address this challenge (Fujimoto et al., 2019; Kumar et al., 2020; Qi et al., 2022; Chen et al., 2023; Zhuang et al., 2023). However, the direct use of current offline RL algorithms in visual control tasks presents two primary difficulties. Initially, offline visual RL is more prone to overfitting issues during representation learning, as it involves extracting hidden states from the limited, high-dimensional visual inputs. Moreover, like its state-space counterpart, offline visual RL is susceptible to the challenge of value overestimation, as we observe from existing methods (Laskin et al., 2020; Hafner et al., 2021).
Figure 1: Our approach solves offline visual RL through a transfer learning paradigm. It harnesses cross-domain knowledge to provide flexible constraints for value estimation on the offline dataset, without impeding state exploration with potential advantages.
Improving offline visual RL remains an under-explored research area. Our goal is to strike a balance between value overestimation and over-conservatism (when excessively penalizing the estimated values beyond the offline data distribution). Intuitively, we should not overly constrain the state exploration with potential advantages. Our basic idea, as illustrated in Figure 1, is to leverage readily available online simulators for related (not necessarily identical) visual control tasks as auxiliary source domains, so that we can frame offline visual RL as an offline-online-offline transfer learning problem to learn mildly conservative policies.
We present a novel model-based transfer RL approach called Collaborative World Models (CoWorld). Specifically, we train separate world models and RL agents for source and target domains, each with domain-specific parameters. To mitigate discrepancies between the world models, we introduce a novel representation learning scheme comprising two iterative training stages. These stages, as shown in Figure 1, facilitate the alignment of latent state distributions (offline to online) and reward functions (online to offline), respectively. By doing so, the source domain critic can serve as an online “test bed” for assessing the target offline policy. It is also more “knowledgeable” as it can actively interact with the online environment and gather rich information. Another benefit of the domain-collaborative world models is the ability to alleviate overfitting issues associated with offline representation learning, leading to more generalizable latent states derived from limited offline visual data.
@@
We showcase the effectiveness CoWorld in offline visual control tasks across the Meta-World, RoboDesk, and DeepMind Control benchmarks. Our approach is shown to be readily extendable to scenarios with multiple source domains. It effectively addresses value overestimation by transferring knowledge from auxiliary domains, even in the presence of diverse physical dynamics, action spaces, reward scales, and visual appearances.
@@
Table 1: Similarities and discrepancies between RoboDesk (target domain) and Meta-World (auxiliary source domain) environments.
@@
Dynamics Simulated Sawyer robot arm Simulated Franka Emika Panda robot arm Different
@@
In summary, our work brings the following contributions:
2 Problem Setup
@@ 3 Method
    Offline-to-online state alignment: Train $\mathcal{M}_{\phi}$ by aligning its state space with that of the source $\mathcal{M}_{\phi^{\prime}}$.
@@
    Online-to-offline value constraint: Train $\{\pi_{\psi},v_{\xi}\}$ with value constraints provided by the source critic $v_{\xi^{\prime}}$.
@@ 3.1 Offline-to-Online State Alignment
We start with a source domain warm-up phase employing a model-based actor-critic method known as DreamerV2 (Hafner et al., 2021). To facilitate cross-domain knowledge transfer, we additionally introduce a state alignment module, which is denoted as $g(\cdot)$ and implemented using the softmax operation. The world model $\mathcal{M}_{\phi^{\prime}}$ consists of the following components:
Recurrent transition: $h_{t}^{(S)}=f_{\phi^{\prime}}(h_{t-1}^{(S)},z_{t-1}^{(S)},a_{t-1}^{(S)})$ (1)
Image encoding: ${e}_{t}^{(S)}=e_{\phi^{\prime}}(o_{t}^{(S)})$
Posterior state: $z_{t}^{(S)}\sim q_{\phi^{\prime}}(h_{t}^{(S)},{e}_{t}^{(S)})$
Prior state: $\hat{z}_{t}^{(S)}\sim p_{\phi^{\prime}}(h_{t}^{(S)})$
Reconstruction: $\hat{o}_{t}^{(S)}\sim p_{\phi^{\prime}}(h_{t}^{(S)},z_{t}^{(S)})$
Reward prediction: $\hat{r}_{t}^{(S)}\sim r_{\phi^{\prime}}(h_{t}^{(S)},z_{t}^{(S)})$
Discount factor: $\hat{\gamma}^{(S)}_{t}\sim p_{\phi^{\prime}}(h_{t}^{(S)},z_{t}^{(S)})$
State alignment target: $s_{t}^{(S)}=g({e}_{t}^{(S)}),$
@@
$\mathcal{L}(\phi^{\prime})=\mathbb{E}_{q_{\phi^{\prime}}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi^{\prime}}(o_{t}^{(S)}\mid h_{t}^{(S)},z_{t}^{(S)})}_{\text{image reconstruction}}$ (2)
$\underbrace{-\ln r_{\phi^{\prime}}(r_{t}^{(S)}\mid h_{t}^{(S)},z_{t}^{(S)})}_{\text{reward prediction}}\underbrace{-\ln p_{\phi^{\prime}}(\gamma_{t}^{(S)}\mid h_{t}^{(S)},z_{t}^{(S)})}_{\text{discount prediction }}$
@@
$\mathcal{L}(\phi)=\mathbb{E}_{q_{\phi}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi}(o_{t}^{(T)}\mid h_{t}^{(T)},z_{t}^{(T)})}_{\text{image reconstruction }}$ (3)
$\underbrace{-\ln r_{\phi}(r_{t}^{(T)}\mid h_{t}^{(T)},z_{t}^{(T)})}_{\text{reward prediction}}\underbrace{-\ln p_{\phi}(\gamma_{t}^{(T)}\mid h_{t}^{(T)},z_{t}^{(T)})}_{\text{discount prediction}}$
$\underbrace{+\ \beta_{1}\mathrm{KL}\left[q_{\phi}(z_{t}^{(T)}\mid h_{t}^{(T)},o_{t}^{(T)})\|\ p_{\phi}(\hat{z}_{t}^{(T)}\mid h_{t}^{(T)})\right]}_{\text{KL divergence}}$
$\underbrace{+\ \beta_{2}\mathrm{KL}\left[\texttt{sg}(g(e_{\phi^{\prime}}(o_{t}^{(T)})))\ \|\ g(e_{\phi}(o_{t}^{(T)}))\right]}_{\text{domain alignment loss}}\Big{]},$
@@
5:     // In the offline domain:
6:     for each step in $\{1:K_{1}\}$ do
7:       Sample $\{(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})\}_{t=1}^{T}\sim\mathcal{B}^{(T)}$.
8:       // Offline-to-online state alignment
9:       Train the target world model $\mathcal{M}_{\phi}$ using Eq. (3).
10:       // Behavior learning with constraint
11:       Generate $\{(z_{i}^{(T)},a_{i}^{(T)})\}_{i=t}^{t+H}$ using $\pi_{\psi}$ and $\mathcal{M}_{\phi}$.
12:       Train the critic $v_{\xi}$ using Eq. (6) over $\{(z_{i}^{(T)},a_{i}^{(T)})\}_{i=t}^{t+H}$.
13:       Train the actor $\pi_{\psi}$ using Eq. (7) over $\{(z_{i}^{(T)},a_{i}^{(T)})\}_{i=t}^{t+H}$.
14:     end for
15:     // In the online domain:
16:     for each step in $\{1:K_{2}\}$ do
17:       Sample $\{(o_{t}^{(S)},a_{t}^{(S)},r_{t}^{(S)})\}_{t=1}^{T}\sim\mathcal{B}^{(S)}$.
18:       // Online-to-offline reward alignment
19:       Sample $\{(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})\}_{t=1}^{T}\sim\mathcal{B}^{(T)}$.
20:       Relabel the source rewards $\{\tilde{r}_{t}^{(S)}\}_{t=1}^{T}$ using Eq. (4).
21:       Train $\mathcal{M}_{\phi^{\prime}}$ using Eq. (2) combined with Eq. (5).
22:       // Source domain behavior learning
23:       Generate $\{(z_{i}^{(S)},a_{i}^{(S)})\}_{i=t}^{t+H}$ using $\pi_{\psi^{\prime}}$ and $\mathcal{M}_{\phi^{\prime}}$.
24:       Train $\pi_{\psi^{\prime}}$ and $v_{\xi^{\prime}}$ over the imagined $\{(z_{i}^{(S)},a_{i}^{(S)})\}_{i=t}^{t+H}$.
25:       Use $\pi_{\psi^{\prime}}$ to collect new source data and append $\mathcal{B}^{(S)}$.
26:     end for
27:  end while
@@ 3.2 Online-to-Offline Reward Alignment
Specifically, we first sample a target domain data trajectory $\{(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})\}_{t=1}^{T}$ from $\mathcal{B}^{(T)}$ (Line 19 in Alg. 1). We then use the source world model parametrized by $\phi^{\prime}$ to extract corresponding latent states and relabel the target-informed source reward (Line 20 in Alg. 1):
$\tilde{h}_{t}=f_{\phi^{\prime}}(\tilde{h}_{t-1},\tilde{z}_{t-1},a_{t-1}^{(T)})$ (4)
$\tilde{e}_{t}=e_{\phi^{\prime}}(o_{t}^{(T)})$
$\tilde{z}_{t}\sim q_{\phi^{\prime}}(\tilde{h}_{t},\tilde{e}_{t})$
$\tilde{r}_{t}^{(S)}=(1-k)\cdot r_{\phi^{\prime}}(\tilde{h}_{t},\tilde{z}_{t})+k\cdot r^{(T)}_{t},$
@@
$\mathcal{L}_{r}(\phi^{\prime})=\ \eta\cdot\mathbb{E}_{\mathcal{B}^{(S)}}\Big{[}\sum_{t=1}^{T}-\ln r_{\phi^{\prime}}(r_{t}^{(S)}|h_{t}^{(S)},z_{t}^{(S)})\Big{]}$ (5)
$+(1-\eta)\mathbb{E}_{\mathcal{B}^{(T)}}\Big{[}\sum_{t=1}^{T}-\ln r_{\phi^{\prime}}(\tilde{r}_{t}^{(S)}|h_{t}^{(T)},z_{t}^{(T)})\Big{]},$
where the second term measures the negative log-likelihood of observing the relabelled source reward $\tilde{r}_{t}^{(S)}$. $\eta$ represents a hyperparameter that gradually decreases from $1$ to $0.1$ throughout this training stage. Intuitively, $\eta$ controls the progressive adaptation of the well-trained source reward predictor to the target domain with limited target reward supervision. We integrate Eq. (5) into Eq. (2) to train the entire world model $\mathcal{M}_{\phi^{\prime}}$ for the source domain agent (Line 21 in Alg. 1) and subsequently perform model-based behavior learning to enable the source critic to assess the target policy (Lines 23-25 in Alg. 1).
@@ 3.3 Min-Max Value Constraint
In the behavior learning phase of the target agent (Lines 11-13 of Alg. 1), we mitigate value overestimation in the offline dataset by introducing a min-max regularization term to the objective function of the target critic model $v_{\xi}$. Initially, we use the auxiliary source critic $v_{\xi^{\prime}}$ to estimate the value function of the imagined target states. Following that, we train $v_{\xi}$ by additionally minimizing the maximum value among the estimates provided by source and target critics:
$\mathcal{L}(\xi)=$ $\ \mathbb{E}_{p_{\phi},p_{\psi}}\Big{[}\sum^{H-1}_{t=1}\underbrace{\frac{1}{2}\left(v_{\xi}(\hat{z}_{t}^{(T)})-\texttt{sg}\big{(}V_{t}^{(T)}\big{)}\right)^{2}}_{\text{value regression}}$ (6)
$+\ \underbrace{\alpha\max\left(v_{\xi}(\hat{z}_{t}^{(T)}),\ \texttt{sg}\big{(}v_{\xi^{\prime}}(\hat{z}_{t}^{(T)})\big{)}\right)}_{\text{value constraint}}\Big{]},$
@@
This approach provides flexibly conservative value estimations, finding a balance between mitigating overestimation and avoiding excessive conservatism in the value function. When the target critic overestimates the value function, the source critic is less vulnerable to the value overestimation problem as it is trained with rich interaction data. Thus, it is possible to observe $v_{\xi}(\hat{z}_{t}^{(T)})>v_{\xi^{\prime}}(\hat{z}_{t}^{(T)})$, and our approach is designed to decrease the output of $v_{\xi}$ to the output of $v_{\xi^{\prime}}$. This prevents the target critic from overestimating the true value. Conversely, when the source critic produces greater values in $v_{\xi^{\prime}}(\hat{z}_{t}^{(T)})$, the min-max regularization term does not contribute to the training of the target critic $v_{\xi}$. This encourages the exploration of potentially advantageous states within the imaginations of the target world model.
In line with DreamerV2 (Hafner et al., 2021), we train the target actor $\pi_{\psi}$ by maximizing a REINFORCE objective function with entropy regularization, allowing the gradients to backpropagate directly through the learned dynamics:
$\mathcal{L}(\psi)$ $=\mathbb{E}_{p_{\phi},p_{\psi}}[\sum_{t=1}^{H-1}(\underbrace{\beta\mathrm{H}[a_{t}^{(T)}\mid\hat{z}_{t}^{(T)}]}_{\text{entropy regularization}}+\underbrace{\rho V_{t}^{(T)}}_{\text{dynamics backprop}}$ (7)
$+\underbrace{(1-\rho)\ln\pi_{\psi}(\hat{a}_{t}^{(T)}\mid\hat{z}_{t}^{(T)})\texttt{sg}(V_{t}^{(T)}-v_{\xi}(\hat{z}_{t}^{(T)})}_{\text{REINFORCE}})].$
@@
Furthermore, it is crucial to note that CoWorld can readily be extended to scenarios with multiple source domains by adaptively selecting a useful task as the auxiliary domain. This extension is easily achieved by measuring the distance of the latent states between the target domain and each source domain. For technical details of the adaptive source domain selection method, please refer to Appendix C.

### v4

@@ Abstract
Training offline RL models using visual inputs poses two significant challenges, i.e., the overfitting problem in representation learning and the overestimation bias for expected future rewards. Recent work has attempted to alleviate the overestimation bias by encouraging conservative behaviors. This paper, in contrast, tries to build more flexible constraints for value estimation without impeding the exploration of potential advantages. The key idea is to leverage off-the-shelf RL simulators, which can be easily interacted with in an online manner, as the “test bed” for offline policies. To enable effective online-to-offline knowledge transfer, we introduce CoWorld, a model-based RL approach that mitigates cross-domain discrepancies in state and reward spaces. Experimental results demonstrate the effectiveness of CoWorld, outperforming existing RL approaches by large margins.
@@ 1 Introduction
Learning control policies with visual observations can be challenging due to high interaction costs with the physical world. Offline reinforcement learning (RL) is a promising approach to address this challenge [11, 21, 37, 3, 65]. However, the direct use of current offline RL algorithms in visual control tasks presents two primary difficulties. Initially, offline visual RL is more prone to overfitting issues during representation learning, as it involves extracting hidden states from the limited, high-dimensional visual inputs. Moreover, like its state-space counterpart, offline visual RL is susceptible to the challenge of value overestimation, as we observe from existing methods [22, 16].
Figure 1: Our approach for offline visual RL.
Improving offline visual RL remains an under-explored research area. We aim to balance between overestimating and over-conservatism of the value function to avoid excessively penalizing the estimated values beyond the offline data distribution. Intuitively, we should not overly constrain the exploration with potential advantages. Our basic idea, as illustrated in Figure 1, is to leverage readily available online simulators for related (not necessarily identical) visual control tasks as auxiliary source domains, so that we can frame offline visual RL as an offline-online-offline transfer learning problem to learn mildly conservative policies.
We present a novel model-based transfer RL approach called Collaborative World Models (CoWorld). Specifically, we train separate world models and RL agents for source and target domains, each with domain-specific parameters. To mitigate discrepancies between the world models, we introduce a novel representation learning scheme comprising two iterative training stages. These stages, as shown in Figure 1, facilitate the alignment of latent state distributions (offline to online) and reward functions (online to offline), respectively. By doing so, the source domain critic can serve as an online “test bed” for assessing the target offline policy. It is also more “knowledgeable” as it can actively interact with the online environment and gather rich information. Another benefit of the domain-collaborative world models is the ability to alleviate overfitting issues of offline representation learning, leading to more generalizable latent states derived from limited offline visual data.
@@
We showcase the effectiveness of CoWorld in offline visual control tasks across the Meta-World, RoboDesk, and DeepMind Control benchmarks. Our approach is shown to be readily extendable to scenarios with multiple source domains. It effectively addresses value overestimation by transferring knowledge from auxiliary domains, even in the presence of diverse physical dynamics, action spaces, reward scales, and visual appearances. In summary, our work brings the following contributions:
@@
2 Problem Setup
Table 1: RoboDesk (target domain) vs. Meta-World (auxiliary source domain).
@@ 2 Problem Setup
Dynamics Simulated Sawyer robot arm Simulated Franka robot arm Different
@@ 3 Method
    Offline-to-online state alignment: Train the offline world model $\mathcal{M}_{\phi}$ by aligning its state space with that of the source world model $\mathcal{M}_{\phi^{\prime}}$.
@@
    Online-to-offline value constraint: Train the target offline-domain agent $\{\pi_{\psi},v_{\xi}\}$ with value constraints provided by the source critic $v_{\xi^{\prime}}$.
@@ 3.1 Offline-to-Online State Alignment
We start with a source domain warm-up phase employing a model-based actor-critic method known as DreamerV2 [16]. To facilitate cross-domain knowledge transfer, we additionally introduce a state alignment module, which is denoted as $g(\cdot)$ and implemented using the softmax operation. The world model $\mathcal{M}_{\phi^{\prime}}$ consists of the following components:
Recurrent transition: $h_{t}^{(S)}=f_{\phi^{\prime}}(h_{t-1}^{(S)},z_{t-1}^{(S)},a_{t-1}^{(S)})$    Image encoding: ${e}_{t}^{(S)}=e_{\phi^{\prime}}(o_{t}^{(S)})$ (1)
Posterior state: $z_{t}^{(S)}\sim q_{\phi^{\prime}}(h_{t}^{(S)},{e}_{t}^{(S)})$    Prior state: $\hat{z}_{t}^{(S)}\sim p_{\phi^{\prime}}(h_{t}^{(S)})$
Reconstruction: $\hat{o}_{t}^{(S)}\sim p_{\phi^{\prime}}(h_{t}^{(S)},z_{t}^{(S)})$    Reward prediction: $\hat{r}_{t}^{(S)}\sim r_{\phi^{\prime}}(h_{t}^{(S)},z_{t}^{(S)})$
Discount factor: $\hat{\gamma}^{(S)}_{t}\sim p_{\phi^{\prime}}(h_{t}^{(S)},z_{t}^{(S)})$    State alignment target: $s_{t}^{(S)}=g({e}_{t}^{(S)}),$
@@
(ϕ)=superscriptitalic-$\mathcal{L}(\phi^{\prime})=$ $\mathbb{E}_{q_{\phi^{\prime}}}\Big{[}\sum_{t=1}^{N}\underbrace{-\ln p_{\phi^{\prime}}(o_{t}^{(S)}\mid h_{t}^{(S)},z_{t}^{(S)})}_{\text{image reconstruction}}\underbrace{-\ln r_{\phi^{\prime}}(r_{t}^{(S)}\mid h_{t}^{(S)},z_{t}^{(S)})}_{\text{reward prediction}}\underbrace{-\ln p_{\phi^{\prime}}(\gamma_{t}^{(S)}\mid h_{t}^{(S)},z_{t}^{(S)})}_{\text{discount prediction }}$ (2)
@@
$\mathcal{L}(\phi)$ $=\mathbb{E}_{q_{\phi}}\Big{[}\sum_{t=1}^{N}\underbrace{-\ln p_{\phi}(o_{t}^{(T)}\mid h_{t}^{(T)},z_{t}^{(T)})}_{\text{image reconstruction }}\underbrace{-\ln r_{\phi}(r_{t}^{(T)}\mid h_{t}^{(T)},z_{t}^{(T)})}_{\text{reward prediction}}\underbrace{-\ln p_{\phi}(\gamma_{t}^{(T)}\mid h_{t}^{(T)},z_{t}^{(T)})}_{\text{discount prediction}}$ (3)
$\underbrace{+\ \beta_{1}\mathrm{KL}\left[q_{\phi}(z_{t}^{(T)}\mid h_{t}^{(T)},o_{t}^{(T)})\|\ p_{\phi}(\hat{z}_{t}^{(T)}\mid h_{t}^{(T)})\right]}_{\text{KL divergence}}\underbrace{+\ \beta_{2}\mathrm{KL}\left[\texttt{sg}(g(e_{\phi^{\prime}}(o_{t}^{(T)})))\ \|\ g(e_{\phi}(o_{t}^{(T)}))\right]}_{\text{domain alignment loss}}\Big{]},$
@@
5:    for each step in $\{1:K_{1}\}$ do $\triangleright$ In the offline domain
6:         Sample $\{(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})\}_{t=1}^{N}\sim\mathcal{B}^{(T)}$.
7:         Train the target world model $\mathcal{M}_{\phi}$ using Eq. (3). $\triangleright$ Offline-to-online state alignment
8:         Generate $\{(z_{i}^{(T)},a_{i}^{(T)})\}_{i=t}^{t+H}$ using $\pi_{\psi}$ and $\mathcal{M}_{\phi}$. $\triangleright$ Behavior learning with constraint
9:         Train the critic $v_{\xi}$ using Eq. (6) over $\{(z_{i}^{(T)},a_{i}^{(T)})\}_{i=t}^{t+H}$.
10:         Train the actor $\pi_{\psi}$ using Eq. (7) over $\{(z_{i}^{(T)},a_{i}^{(T)})\}_{i=t}^{t+H}$.
11:    end for
12:    for each step in $\{1:K_{2}\}$ do $\triangleright$ In the online domain
13:         Sample $\{(o_{t}^{(S)},a_{t}^{(S)},r_{t}^{(S)})\}_{t=1}^{N}\sim\mathcal{B}^{(S)}$.
14:         Sample $\{(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})\}_{t=1}^{N}\sim\mathcal{B}^{(T)}$. $\triangleright$ Online-to-offline reward alignment
15:         Relabel the source rewards $\{\tilde{r}_{t}^{(S)}\}_{t=1}^{N}$ using Eq. (LABEL:eq:corrected_reward).
16:         Train $\mathcal{M}_{\phi^{\prime}}$ using Eq. (2) combined with Eq. (5).
17:         Generate $\{(z_{i}^{(S)},a_{i}^{(S)})\}_{i=t}^{t+H}$ using $\pi_{\psi^{\prime}}$ and $\mathcal{M}_{\phi^{\prime}}$. $\triangleright$ Source domain behavior learning
18:         Train $\pi_{\psi^{\prime}}$ and $v_{\xi^{\prime}}$ over the imagined $\{(z_{i}^{(S)},a_{i}^{(S)})\}_{i=t}^{t+H}$.
19:         Use $\pi_{\psi^{\prime}}$ to collect new source data and append $\mathcal{B}^{(S)}$.
20:    end for
21:end while
@@ 3.2 Online-to-Offline Reward Alignment
Specifically, we first sample a target domain data trajectory $\{(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})\}_{t=1}^{T}$ from $\mathcal{B}^{(T)}$ (Line 14 in Alg. 1). We then use the source world model parametrized by $\phi^{\prime}$ to extract corresponding latent states and relabel the target-informed source reward (Line 15 in Alg. 1):
$\tilde{h}_{t}=f_{\phi^{\prime}}(\tilde{h}_{t-1},\tilde{z}_{t-1},a_{t-1}^{(T)})$ $\tilde{e}_{t}=e_{\phi^{\prime}}(o_{t}^{(T)})$ (4)
$\tilde{z}_{t}\sim q_{\phi^{\prime}}(\tilde{h}_{t},\tilde{e}_{t})$ $\tilde{r}_{t}^{(S)}=(1-k)\cdot r_{\phi^{\prime}}(\tilde{h}_{t},\tilde{z}_{t})+k\cdot r^{(T)}_{t},$
@@
$\mathcal{L}_{r}(\phi^{\prime})=\ \eta\cdot\mathbb{E}_{\mathcal{B}^{(S)}}\Big{[}\sum_{t=1}^{N}-\ln r_{\phi^{\prime}}(r_{t}^{(S)}|h_{t}^{(S)},z_{t}^{(S)})\Big{]}+(1-\eta)\mathbb{E}_{\mathcal{B}^{(T)}}\Big{[}\sum_{t=1}^{N}-\ln r_{\phi^{\prime}}(\tilde{r}_{t}^{(S)}|h_{t}^{(T)},z_{t}^{(T)})\Big{]},$ (5)
where the second term measures the negative log-likelihood of observing the relabelled source reward $\tilde{r}_{t}^{(S)}$. $\eta$ represents a hyperparameter that gradually decreases from $1$ to $0.1$ throughout this training stage. Intuitively, $\eta$ controls the progressive adaptation of the well-trained source reward predictor to the target domain with limited target reward supervision. We integrate Eq. (5) into Eq. (2) to train the entire world model $\mathcal{M}_{\phi^{\prime}}$ for the source domain agent (Line 16 in Alg. 1) and subsequently perform behavior learning to enable the source critic to assess the target policy (Lines 17-19 in Alg. 1).
@@ 3.3 Min-Max Value Constraint
In the behavior learning phase of the target agent (Lines 8-10 of Alg. 1), we mitigate value overestimation in the offline dataset by introducing a min-max regularization term to the objective function of the target critic model $v_{\xi}$. Initially, we use the auxiliary source critic $v_{\xi^{\prime}}$ to estimate the value function of the imagined target states. Following that, we train $v_{\xi}$ by additionally minimizing the maximum value among the estimates provided by source and target critics:
$\mathcal{L}(\xi)=\ \mathbb{E}_{p_{\phi},p_{\psi}}\Big{[}\sum^{H-1}_{t=1}\underbrace{\frac{1}{2}\left(v_{\xi}(\hat{z}_{t}^{(T)})-\texttt{sg}\big{(}V_{t}^{(T)}\big{)}\right)^{2}}_{\text{value regression}}+\ \underbrace{\alpha\max\left(v_{\xi}(\hat{z}_{t}^{(T)}),\ \texttt{sg}\big{(}v_{\xi^{\prime}}(\hat{z}_{t}^{(T)})\big{)}\right)}_{\text{value constraint}}\Big{]},$ (6)
@@
This approach provides flexibly conservative value estimations, finding a balance between mitigating overestimation and avoiding excessive conservatism in the value function. When the target critic overestimates the value function, the source critic is less vulnerable to the value overestimation problem as it is trained with rich interaction data. Thus, it is possible to observe $v_{\xi}(\hat{z}_{t}^{(T)})>v_{\xi^{\prime}}(\hat{z}_{t}^{(T)})$, and our approach is designed to decrease the output of $v_{\xi}$ to the output of $v_{\xi^{\prime}}$. This prevents the target critic from overestimating the true value. Conversely, when the source critic produces greater values in $v_{\xi^{\prime}}(\hat{z}_{t}^{(T)})$, the min-max regularization term does not contribute to the training of the target critic $v_{\xi}$. This encourages the exploration of potentially advantageous states within the imaginations of the target world model. In line with DreamerV2 [16], we train the target actor $\pi_{\psi}$ by maximizing a REINFORCE objective function with entropy regularization, allowing the gradients to backpropagate directly through the learned dynamics:
$\mathcal{L}(\psi)=\mathbb{E}_{p_{\phi},p_{\psi}}\sum_{t=1}^{H-1}(\underbrace{\beta\mathrm{H}[a_{t}^{(T)}\mid\hat{z}_{t}^{(T)}]}_{\text{entropy regularization}}+\underbrace{\rho V_{t}^{(T)}}_{\text{dynamics backprop}}+\underbrace{(1-\rho)\ln\pi_{\psi}(\hat{a}_{t}^{(T)}\mid\hat{z}_{t}^{(T)})\texttt{sg}(V_{t}^{(T)}-v_{\xi}(\hat{z}_{t}^{(T)})}_{\text{REINFORCE}}).$ (7)
@@
Furthermore, it is crucial to note that CoWorld can readily be extended to scenarios with multiple source domains by adaptively selecting a useful task as the auxiliary domain. This extension is easily achieved by measuring the distance of the latent states between the target domain and each source domain. For technical details of the adaptive source domain selection, please refer to Appendix C.